OpenStudy (anonymous):

how to find dw/dx of w= x+4

OpenStudy (jhannybean):

Just take the derivative with respect to x,

OpenStudy (anonymous):

HOw?

OpenStudy (anonymous):

the answer is 1 but i cant understand how?

OpenStudy (jhannybean):

\[w=x+4\]\[\frac{dw}{dx} = \frac{dx}{dx}+\frac{d4}{dx}\]This is just like taking the derivative. Have you learned how to take derivatives?

OpenStudy (anonymous):

no

OpenStudy (jhannybean):

Because w is a FUNCTION of x, when we're taking the derivative of a variable with respect to another variable, we label that as a prime, \(w'\). If we're taking the derivative of a variable with respect to itself, it depends upon the power rule of the variable. in this case, \(\frac{d}{dx}(x^1) = 1\)

OpenStudy (jhannybean):

These are just rules in your Calculus book you have to follow. http://www.mathsisfun.com/calculus/derivatives-introduction.html use this for more information about finding derivatives.

OpenStudy (anonymous):

@Ghayyas.mubashir Do you know the definition of a derivative?

OpenStudy (anonymous):

Have you seen this: \[ \frac{df}{dx} = \lim_{h\to 0}\frac{f(x+h)-f(x)}{h} \]

OpenStudy (anonymous):

yeh but i cant understand that

OpenStudy (anonymous):

Okay, let's start from the basics then...

OpenStudy (anonymous):

yes plzz

OpenStudy (anonymous):

\[ w = f(x) = x+4 \]Do you understand this?

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

Okay, now: \[ f(\color{red}{x+h}) = (\color{red}{x+h})+4 \]Does this make sense?

OpenStudy (anonymous):

We substituted in \(x+h\) for \(x\).

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

Now we plug these into the limit. \[ \lim_{h\to 0}\frac{f(x+h)-f(x)}{h}=\lim_{h\to 0}\frac{(x+h+4)-(x+4)}{h} \]

OpenStudy (anonymous):

Do you understand the substitutions I did here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can you simplify: \[ \lim_{h\to 0}\frac{(x+h+4)-(x+4)}{h} \]?

OpenStudy (anonymous):

wait a min i try

OpenStudy (anonymous):

i cant

OpenStudy (anonymous):

\[ \lim_{h\to 0}\frac{(x+h+4)-(x+4)}{h} = \lim_{h\to 0}\frac{x+h+4-x-4}{h} =\lim_{h\to 0}\frac{h}{h}=\lim_{h\to0}1=1 \]?

OpenStudy (anonymous):

\[ \frac{dw}{dx}=\frac{df}{dx}=1 \]

OpenStudy (anonymous):

May god help you if you ever are tested on this.

OpenStudy (anonymous):

h/h =0 then y is equal to 1?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

You make me sad.

OpenStudy (anonymous):

soory but plz if u can

OpenStudy (anonymous):

\[ \frac 44 = ? \]

OpenStudy (anonymous):

1

OpenStudy (anonymous):

\[ \frac hh = ? \]

OpenStudy (anonymous):

but h = 0

OpenStudy (anonymous):

0/0 ? = 1?

OpenStudy (anonymous):

No, that isn't how limits work.

OpenStudy (anonymous):

any number that has limit 0 it would be always one??

OpenStudy (anonymous):

When you are doing a limit, like \[ \lim_{x\to a}f(x) \]Then you assume that \(x\) is very close to \(a\), but that \(x\neq a\).

OpenStudy (dan815):

hey ghayyas

OpenStudy (jhannybean):

You should really use the link I provided you in the posts above to learn about limits and derivatives to get a general idea of how they work.

OpenStudy (dan815):

you are learning first principle?

OpenStudy (dan815):

first principle is basically comes from the equation of a line

OpenStudy (anonymous):

hi dan815 abd thank you much jhanny bean and best regards wio

OpenStudy (anonymous):

dian i understand now. thank you so much

OpenStudy (dan815):

okay

OpenStudy (dan815):

np

OpenStudy (anonymous):

a big thans to wio who help me alot.. Thank you SOOOOO MUCH,,

OpenStudy (anonymous):

You're welcome. "First principle" Where does that even come from?

OpenStudy (anonymous):

Is there a second principle?

OpenStudy (anonymous):

what is first principle>?

OpenStudy (jhannybean):

Yeah that's what i'm wondering.